# Non-Hermitian ultra-strong bosonic condensation through interaction-induced caging

### v1

@@ Abstract
Introduction.— In many-body physics, interactions between particles play a crucial role, driving interesting phenomena such as phase transitions [1, 2, 3, 4, 5], superconductivity [6, 7, 8], and quantum entanglement [9, 10, 11, 5]. Much of these physics can be observed through many-body dynamical simulations in both classical and quantum simulators [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28].
There has been growing excitement surrounding the interplay between many-body interactions and the non-Hermitian skin effect (NHSE) [29, 30, 31, 32, 33, 34, 35]. This excitement stems from the non-locality and sensitivity of the NHSE, which has a profound impact on particle localization [36, 37, 38, 39, 40, 41, 42] and condensation dynamics [43, 44, 45, 46, 42, 47], which can result in unexpected consequences when interplayed with quantum many-body interactions [29, 30, 31, 32, 33, 34, 35]. These include the emergence of a Fermi surface in a many-body non-Hermitian fermionic chain [29], exotic spin liquids from couplings to the environment [30], symmetry breaking and spectral structure of the interacting Hatano-Nelson model [31], and unconventional emission and absorption rates in non-Hermitian baths, as well as fundamental modifications of the quantum dynamics [32, 34]. Recently, the non-Hermitian Mott skin effect [48] has revealed the intriguing interplay between strong correlations and non-Hermitian topology in the many-body skin effect.
In this work, we uncovered a new mechanism whereby bosons can be made to condense very strongly through the triple interplay of non-Hermitian pumping, density interactions and non-trivial band topology. Due to the emergent non-locality of both the NHSE and the density interactions, the observed phenomenon is in stark contrast to the conventional behavior of interacting bosons, such as photons, which repel each other when they interact, leading to the quantum phenomenon known as photon blockade [49, 50, 51, 52]. Our simulations show that ultra-strong localization, and not just boundary state accumulation, is achieved when both non-Hermiticity and density interactions are present, which is not attainable by either alone.
Bosonic condensation from interactions and non-Hermiticity.– To demonstrate how non-Hermitian pumping, topology and bosonic interactions can interplay to cause unexpectedly strong bosonic condensation, we consider a minimal interacting 1D bosonic lattice model that contains topological edge modes at its boundary sites:
@@
$+\sum_{x=1}^{L-1}\left(b_{2x}^{\dagger}b_{2x+1}+b_{2x+1}^{\dagger}b_{2x}\right)+\frac{U}{2}n_{x_{0}}^{2},$
@@
which is a Su-Schrieffer-Heeger (SSH) lattice [53] with asymmetric intercell ($t_{L}$ and $t_{R}$) hoppings and unit intracell hoppings, equipped with a density interaction $\frac{1}{2}Un_{x_{0}}^{2}$ at site $x_{0}$, $U\geq 0$. Here $b_{x}$ ($b_{x}^{\dagger}$) is the bosonic annihilation (creation) operator at sites $x=1,2,\ldots,2L$, with $n_{x_{0}}=b_{x_{0}}^{\dagger}b_{x_{0}}$ the boson number operator. Non-Hermitian pumping [36, 54, 37, 55, 38, 56, 42, 57, 58, 59, 60, 44, 61, 40, 62, 63, 64, 65, 41, 66, 67, 68, 69, 70, 71, 72, 73, 47] occurs towards the left if $r=t_{L}/t_{R}>1$, accumulating the bosons towards the left boundary. However, when there are multiple indistinguishable bosons, this accumulation would also be affected by the repulsive ($U>0$) interaction at site $x_{0}$, with strength depending on bosonic occupancy. In this work, we have chosen the interaction $\frac{1}{2}Un_{x_{0}}^{2}$ that mimics the simplest possible nonlinearity in the mean-field limit, different from the usual Bose-Hubbard interaction [15, 49, 74, 50, 51, 52] by a local density shift [75, 76, 77, 78, 79].
@@
Figure 1: Spatial density accumulation (a-d) and ultra-strong bosonic condensation (e) due to a boundary density interaction at $x_{0}=1$, for different two-boson initial states (1,1), (4,4), and (1,4) evolving via $H$ [Eq. 1]. $r=1$ or $4$ corresponds to Hermitian (a,b) or non-Hermitian (c,d) cases, and $U=0$ or $U\neq 0$ corresponds the absence (a,c) or presence (b,d) of the density interaction. (a-d) Heatmaps depict $\rho_{x}(t)$ [Eq. 3], the time-dependent boson number over the simulation duration $t\in[0,T]$ with $T=10$. The line plot above each heatmap shows the corresponding time-averaged boson number profile $\bar{\rho}_{i}$ [Eq. 4]. The bosons are most likely to be trapped at site $x=1$ in the presence of both leftwards NHSE ($r>0$) and nonzero interaction ($U>0$), as shown by the white and bright yellow regions in the heatmaps. While normally being considered as repulsive, the interaction term $U$ also acts as a trap at $x_{0}=1$ in non-equilibrium dynamics. (e) The dramatic interplay of NHSE pumping and bosonic interaction is most saliently expressed through the bosonic condensation at $x=1$, as given by the two-boson correlation $P_{11}$ [Eq. 5]. Here, we observe significantly larger $P_{11}$, indicated by the red curves in all three (e) subfigures, which demonstrates the greatly enhanced ability to trap both bosons at site 1 when both $U>0$ and $r>0$.
@@
While one might naively expect a repulsive ($U>0$) density interaction to suppress the NHSE accumulation primarily, the observed behavior can be dramatically different even with just two bosons. Two distinct measures of particle accumulation can be defined: the spatial density accumulation and the many-body correlation. An initial 2-boson state $|\phi(t=0)\rangle$ evolves according to the Schrodinger equation $|\dot{\phi}(t)\rangle=-iH|\phi(t)\rangle$, and can be expressed as
@@
Figure 2: Extent of two-boson condensation due to a boundary density interaction at $x_{0}=1$ [Eq. 1], and its correspondence with the boundary localization of 2-boson spectral bands. (a) Time-averaged two-boson condensation probability $\bar{P}_{11}$ [Eq. 6] at site $x=1$, in the parameter space of non-Hermitian hopping asymmetry $r$ and density interaction strength $U$. (b) Corresponding two-boson spectra at $r=4$, which features five bands, with bands 2 and 4 being containing a pair of interaction-hybridized bulk and topological bosons. As $U$ increases, it creates a group of eigenstates at $E\approx U$ that exhibits strong localization $\psi_{11}$ (red) at $x=x_{0}=1$. Strong $\psi_{11}$ localization in the hybrid topological band 4 leads to suppressed $\bar{P}_{11}$ for the boundary-localized initial state $(1,1)$ (circled in dashed black), but enhanced $\bar{P}_{11}$ condensation for the initial state $(1,4)$ (solid black). By contrast, strong $\psi_{11}$ localization in the bulk bands 3 and 5 (circled in light blue) corresponds to enhanced $\bar{P}_{11}$ for the bulk initial state $(4,4)$.
@@
What is striking, however, is the ultra-strong two-boson condensation at site $x_{0}=1$ when both non-Hermiticity ($r>1$) and the density interaction ($U>0$) are present. This is revealed in the dynamical $P_{11}(t)$ [Eq. 5] plots of Fig. 1(e), which shows the probability of both bosons condensing at $x=1$. From the Top Row, when both bosons are prepared at $x=1$ (initial state =(1,1)), they can only remain both there if non-Hermiticity and interaction are simultaneously present (red). But most surprising is what happens when at least one boson is initially away from $x=1$ (Center and Bottom Rows): we observe overwhelmingly higher $P_{11}$ only when non-Hermiticity and interaction simultaneously interplay (red), compared to having either on their own (yellow, green). This suggests that the ability to trap both bosons and suppress photon blockade [50, 51, 49, 52] hinges on an emergent consequence of this interplay, a fact not evident from the density evolution alone [Fig. 1(a-d)].
@@
Figure 3: Caging mechanism of interaction term $U$ with initial state (2,2), left of interaction position $x_{0}=3$. (a) Snapshots of smoothed time-averaged two-boson correlation probability $\bar{P}_{11}$ (Eq. 6) at $t=0,2.5,5,8$ for (a1) NH non-interacting and (a2) NH interacting cases. An inverted L-shape indicates an effective interaction range. (b1) Probability of both bosons at site 1. NH interacting case (red) shows optimal trapping. (b2) Single-boson probability ($\bar{P}_{1}$) at site 1 for comparison. Surprisingly, the interaction term acts as a barrier, confining bosons to sites 1-3 in the NH interacting case, demonstrating a unique caging effect not observed in single-boson scenarios or non-interacting cases. This highlights the distinctive role of interactions in two-boson dynamics. Lattice: $L=5$ cells; $U=1$; hopping asymmetry $r=4$; smoothing $\Delta t=2$.
Figure 4: Caging mechanism of interaction term $U$ with initial state (4,4), right of interaction position $x_{0}=3$. (a) Snapshots of smoothed time-averaged two-boson correlation probability $\bar{P}_{11}$ (Eq. 6) at $t=0,1,1.6,2.2,3,8$ for (a1) NH non-interacting and (a2) NH interacting cases. Initial similar behavior in both cases ($t=0,1$). Interaction barrier effect emerges ($t=1.6,2.2$). NHSE-assisted confinement to left side in interacting case ($t=3,8$). (b1) Smoothed probability of both bosons at site 1 ($\bar{P}_{11}$ in Eq. 6). (b2) Single-boson probability ($\bar{P}_{1}$) for comparison. Notably, the NH interacting case demonstrates superior trapping ability for two bosons compared to the single-boson scenario, highlighting a unique interplay between interaction and non-Hermiticity. $U=6$ and smoothing $\Delta t=2$. (c) Time-averaged two-boson probability $\bar{P}{mn}$ (Eq. 6) within the cage, $\sum{m,n=1}^{x_{0}}\bar{P}_{mn}$, plotted against hopping asymmetry values $\log r\in[0,6]$ for various initial states: (c1) (4,4), (c2) (2,2), and (c3) (2,4). The orange curves ($U=1$) in (c1) and (c2) confirm the caging mechanism. A slight deviation for the orange curve ($U=1$) compared with the blue curve ($U=0$) is observed in (c3), attributed to insufficient interaction time between the two bosons. Lattice size $L=5$ unit cells.
@@
Topological origin of ultra-strong bosonic condensation.– Intriguingly, for a density interaction at the $x_{0}=1$ boundary, it turns out that the ultra-strong boundary bosonic condensation is not just due to the interaction-NHSE interplay, but also relies crucially on SSH topology. This is evident from the 2-boson band structure plot in Fig. 2(b), where each eigenenergy $E$ is colored according to the overlap $\psi_{11}=\langle(1,1)|\psi\rangle$ of its corresponding eigenstate $|\psi\rangle$ with $|(1,1)\rangle=(b_{1}^{\dagger})^{2}|0\rangle$. Computed for our Hamiltonian $H$ [Eq. 1] at strong non-Hermitian hopping asymmetry $r=4$, it features 5 bands, with bands 2 and 4 arising from the hybridization of bulk and topological bands. This follows from the single-boson band structure [80], which contains two symmetrically gapped bulk bands separated by in-gap topological zero modes. The effect of the $U$ density interaction is to induce high $\psi_{11}$ overlap (red) successively from bands 3 to 5, as $U$ is increased.
@@
To illustrate, with the initial state (1,1), suppressed $\bar{P}_{11}$at $U\approx 3$ corresponds to high $\psi_{11}$ overlap with band 4 [circled in black in Figs.2(a,b)]. This is because band 4 manifestly contain one bulk and one topological boson, which is not consistent with maintaining both bosons at $x_{0}=1$. As such, the same interaction strength of $U\approx 3$ leads to enhanced $\bar{P}_{11}$ (dashed black) for the initial state (1,4), where only one boson overlaps with the topological boundary mode. Likewise, $U\approx 1$ or $U\approx 5$ favors the $\bar{P}_{11}$ bosonic condensation for the bulk initial state (4,4), as circled in light blue, since the $\psi_{11}$ overlap with the bulk bands (3 and 5) are strongest at these interaction strengths $U$.
@@
2-boson interactions as caging mechanisms.– Interestingly, the boson density interaction can also act as a non-local cage when the interaction position $x_{0}$ is not at the boundary site $1$. To demonstrate this, we shall show in Fig. 3 that, for an initial state with both bosons to the left of $x_{0}$, an interaction $U$ cages the evolved state, trapping it. And, more elaborately, we show in Fig. 4 that, for an initial state with both bosons to the right of $x_{0}$, the interaction will first act as a barrier to the evolved state, but will eventually still end up as a cage that traps the particle flux that leaked into it via NHSE pumping. We shall consider $x_{0}=3$ and initial states $(2,2)$ and $(4,4)$ in Fig. 3 and Fig. 4 respectively.
@@
To illustrate the caging mechanism more intuitively, we express our 1D bosonic interacting model in its 2D configuration space $(x,x^{\prime})$, where $x$ and $x^{\prime}$ represent the positions of the two bosons. Here $(x,x^{\prime})$ and $(x^{\prime},x)$ refer to the same state due to bosonic statistics [74, 81, 82, 83, 77, 84]. A key observation is that the nonlinear term $\frac{1}{2}Un_{x_{0}}^{2}$ can be interpreted as a non-local potential offset [74, 77].
When the initial state is already inside the cage, bosons are predominantly confined within the cage (inverted-L shape) between sites $1$ and $3$, as shown in Fig. 4(a2). This contrasts with Fig. 4(a1), which shows no such suppression of spreading, despite the pumping direction being towards the left-top corner. When the initial state is outside the cage, as in Fig. 4(a), bosons are pumped to the left side of the lattice due to skin pumping. However, once they enter the cage, they become trapped. This is clearly demonstrated in Fig. 4(a2) at $t=3$ and $t=8$, where bosons are predominantly confined to the left side of the lattice, explaining the prominence of the red curve in Fig. 4(b1).
To prove that this caging mechanism is unique to the interacting scenario and not merely a consequence of an effective on-site potential, we conducted a comparative analysis. We examined the probability ($\bar{P}_{1}$, smoothed using the same method as in the two-boson case) of a single boson localized at the edge (site $1$) of a 1D SSH chain with an on-site potential $U$ at site $1$. The single-boson case is not as effective as the two-boson case, as shown in Fig. 3(b2) and Fig. 4(b2), which stand in stark contrast to the outstanding red curves in Fig. 3(b1) and Fig. 4(b1). This further proves that the interaction effect is unique to the two-boson case.
The efficacy of the caging mechanism is corroborated by analyzing the time-averaged two-boson probability within the caging region, $\sum_{m,n=1}^{x_{0}}\bar{P}_{mn}$, as a function of hopping asymmetry $\log r$ for diverse initial states (Fig.4(c)) in long-time dynamics over a period of 500. In Fig.4(c1) and (c2), the significantly higher values of the interacting case ($U=1$, orange curves) compared to the non-interacting case ($U=0$, blue curves) provide strong evidence for the caging phenomenon. Notably, Fig. 4(c3) exhibits a small deviation between the interacting and non-interacting scenarios, which we attribute to limited interaction time between the bosons when they start on opposite sides of the interaction site. This leads to a less pronounced caging effect, as the boson starting on site 2 is less likely to occupy site 3 due to the leftward skin pumping. This observation also explains the small difference between orange and blue curves in Fig. 3(b2) and Fig. 4(b2) – the caging effect is significant only when two bosons are interacting.
@@
Conclusion.— To summarize, our investigation has revealed how the intriguing interplay between NHSE, topology and boson-boson interactions can unexpectedly lead to ultra-strong edge condensation of the bosons. The extent of this particle condensation goes beyond intuitive expectations of NHSE-pumped bosons trapped by the interaction, being also facilitated by the interaction-induced hybridization of topological and bulk states. Moreover, even more pronounced bosonic condensation can be induced by density interactions in the bulk ($x_{0}\neq 1$) due to an emergent caging mechanism. This trapping phenomenon is not observed in single-boson dynamics, and generalizes to higher numbers of bosons as well as generic density interactions that are non-local in the many-body configuration space. Our findings provide new perspectives on the role of interactions in the condensation and localization of bosons in non-Hermitian topological systems, offering a novel approach to trapping and controlling bosons. Our results are not limited to the specific model we have studied but can be extended to other systems with different forms of interactions and topologies. Experimentally, various platforms can potentially implement the setup for locally interacting bosons, including optical lattices [85, 86, 87, 18, 88, 89], photonic systems [90, 91, 92, 93, 94, 95, 75], and circuit QED [96, 25, 97, 98, 99]. Specifically, we have a potential experimental realization of the bosonic interacting model using resonator arrays in real space in the Supplemental Material [80].
Supplemental Material for
@@
“Non-Hermitian ultra-strong bosonic condensation through interaction-induced caging”

### v2

@@ Abstract
Introduction.— There has been growing excitement in interplay between many-body interactions and the non-Hermitian skin effect (NHSE) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. This stems from the non-locality and sensitivity of the NHSE [22, 23, 14, 24, 25, 26, 3, 27, 28, 5], which profoundly impacts particle localization [1, 2, 3, 4, 5, 6, 7] and condensation dynamics [8, 29, 9, 10, 11, 12], which leads to unexpected consequences in the presence of quantum many-body interactions. These include the emergence of a many-fermion Fermi skin [2, 7], exotic spin liquids from environmental couplings [1], spectral symmetry breaking [13], unconventional quantum dynamics in non-Hermitian baths [9, 30, 10, 31, 32, 33] and the non-Hermitian Mott skin effect [20].
In this work, we uncovered a new mechanism whereby bosons can be made to condense very strongly through the triple interplay of non-Hermitian pumping, density interactions and non-trivial band topology. Due to the emergent non-locality of both the NHSE [22, 23, 14, 24, 25] and the density interactions, the observed phenomenon contrasts starkly with the conventional behavior of interacting bosons, such as photons, which repel each other when they interact i.e. photon blockade [34, 35, 36, 37]. In particular, we show that ultra-strong localization, and not just boundary state accumulation, arises from an emergent caging mechanism that requires both non-Hermitian pumping and density interactions.
Ultra-strong non-Hermitian Bosonic condensation.— To demonstrate how non-Hermitian pumping, topology and bosonic interactions can interplay to cause unexpectedly strong bosonic condensation, we consider a minimal interacting 1D bosonic lattice model that contains topological edge modes at its boundary sites:
@@
$+\sum_{x=1}^{L-1}t_{0}\left(b_{2x}^{\dagger}b_{2x+1}+b_{2x+1}^{\dagger}b_{2x}\right)+\frac{U}{2}n_{x_{0}}^{2},$
@@
which is a Su-Schrieffer-Heeger (SSH) lattice [38] with asymmetric intercell ($t_{L}$ and $t_{R}$) hoppings and unit intracell hoppings, equipped with a density interaction $\frac{1}{2}Un_{x_{0}}^{2}$ at site $x_{0}$, $U\geq 0$. Here $b_{x}$ ($b_{x}^{\dagger}$) is the bosonic annihilation (creation) operator at sites $x=1,2,\ldots,2L$, with $n_{x_{0}}=b_{x_{0}}^{\dagger}b_{x_{0}}$ the boson number operator. Non-Hermitian pumping [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 24, 50, 51, 52, 53, 54, 22, 55, 56, 57, 58, 59, 60, 61, 14, 62, 63, 64, 65, 66, 67, 68, 69, 70] occurs towards the left if $r=t_{L}/t_{R}>1$, accumulating the bosons towards the left boundary. However, when there are multiple indistinguishable bosons, this accumulation would also be affected by the repulsive ($U>0$) interaction at site $x_{0}$, with strength depending on bosonic occupancy. In this work, we have chosen the interaction $\frac{1}{2}Un_{x_{0}}^{2}$ that mimics the simplest possible nonlinearity in the mean-field limit, different from the usual Bose-Hubbard interaction [71, 34, 72, 35, 36, 37] by a local density shift [73, 74, 75, 76, 77].
@@
Figure 1: Spatial density accumulation (a-d) and ultra-strong bosonic condensation (e) due to a boundary density interaction at $x_{0}=1$, for different two-boson initial states (1,1), (4,4), and (1,4) evolving via $H$ [Eq. 1]. $r=1$ or $4$ corresponds to Hermitian (a,b) or non-Hermitian (c,d) cases, and $U=0$ or $U\neq 0$ corresponds the absence (a,c) or presence (b,d) of the density interaction. (a-d) Heatmaps depict $\rho_{x}(t)$ [Eq. 3], the time-dependent boson number over the simulation duration $t\in[0,T]$ with $T=10$. The line plot above each heatmap shows the corresponding time-averaged boson number profile $\bar{\rho}_{i}$ [Eq. 4]. The bosons are most likely to be trapped at site $x=1$ in the presence of both leftwards NHSE ($r>0$) and nonzero interaction ($U>0$), as shown by the white and bright yellow regions in the heatmaps. While normally being considered as repulsive, the interaction term $U$ also acts as a trap at $x_{0}=1$ in non-equilibrium dynamics. (e) The dramatic interplay of NHSE pumping and bosonic interaction is most saliently expressed through the bosonic condensation at $x=1$, as given by the two-boson correlation $P_{11}$ [Eq. 5]. Here, we observe significantly larger $P_{11}$, indicated by the red curves in all three (e) subfigures, which demonstrates the greatly enhanced ability to trap both bosons at site 1 when both $U>0$ and $r>0$. $t_{0}=3$ for all cases. $r=4$ corresponds to $t_{L}=1.6$ and $t_{R}=0.4$.
@@
While one might naively expect a repulsive ($U>0$) density interaction to primarily suppress the NHSE, the observed behavior can be dramatically different even with just two bosons. Two distinct measures of particle accumulation can be defined: the spatial density accumulation and the many-body correlation. An initial 2-boson state $|\phi(t=0)\rangle$ evolves according to the Schrodinger equation $|\dot{\phi}(t)\rangle=-iH|\phi(t)\rangle$, and can be expressed as
@@
What is striking, however, is the ultra-strong two-boson condensation at boundary site $x_{0}=1$ when non-Hermiticity ($r>1$) and the density interaction ($U>0$) are both present. This is revealed in the dynamical $P_{11}(t)$ [Eq. 5] plots in Fig. 1(e), which shows the probability of both bosons condensing at $x=1$. From the Top Row, when both bosons are prepared at $x=1$ (initial state =(1,1)), they can only remain both there if the interaction and non-Hermiticity are simultaneously present (red). But most surprising is what happens when at least one boson is initially away from $x=1$ (Center and Bottom Rows): we observe overwhelmingly higher $P_{11}$ only when non-Hermiticity and interaction simultaneously interplay (red), compared to having either on their own (yellow, green). This suggests that trapping both bosons and suppress photon blockade [35, 36, 34, 37] hinges on an emergent consequence of this interplay, a fact not evident from the density evolution alone [Fig. 1(a-d)].
@@
Figure 2: Extent of two-boson condensation due to a boundary density interaction at $x_{0}=1$ [Eq. 1], and its correspondence with the boundary localization of 2-boson spectral bands. (a) Time-averaged two-boson condensation probability $\bar{P}_{11}$ [Eq. 6] at site $x=1$, in the parameter space of non-Hermitian hopping asymmetry $r$ and density interaction strength $U$. (b) Corresponding two-boson spectra at $r=4(t_{L}=1.6$ and $t_{R}=0.4),t_{0}=3$, which features five bands, with bands 2 and 4 being containing a pair of interaction-hybridized bulk and topological bosons. As $U$ increases, it creates a group of eigenstates at $E\approx U$ that exhibits strong localization $\psi_{11}$ (red) at $x=x_{0}=1$. Strong $\psi_{11}$ localization in the hybrid topological band 4 leads to suppressed $\bar{P}_{11}$ for the boundary-localized initial state $(1,1)$ (circled in dashed black), but enhanced $\bar{P}_{11}$ condensation for the initial state $(1,4)$ (solid black). By contrast, strong $\psi_{11}$ localization in the bulk bands 3 and 5 (circled in light blue) corresponds to enhanced $\bar{P}_{11}$ for the bulk initial state $(4,4)$.
Figure 3: Ultrastrong boundary condensation $\bar{P}_{11}$ from interaction-induced caging, for initial state (2,2) within a $x_{0}=3$ cage for a 5 unit cell chain. (a) Snapshots of the time-smoothed ($\Delta t=2$) two-boson correlation probability $\bar{P}_{11}$ [Eq. 6] for (a1) non-interacting and (a2) interacting cases. The L-shaped cage in (a2) traps the bosons tightly. (b1) Evolution of probability of both bosons at the left boundary, with greatly enhanced $\bar{P}_{11}(t)$ only for the non-Hermitian interacting case (red). (b2) When only one boson is present, the boundary density $\bar{\rho}_{1}(t)$ remains low regardless of whether a $U$ barrier is present, showcasing that the caging mechanism is an interaction effect.
@@
Topological origin of ultra-strong bosonic condensation.— Intriguingly, for a density interaction at the $x_{0}=1$ boundary, it turns out that the ultra-strong boundary bosonic condensation is not just due to the interaction-NHSE interplay, but also relies crucially on SSH topology. This is evident from the 2-boson band structure plot in Fig. 2(b), where each eigenenergy $E$ is colored according to the overlap $\psi_{11}=\langle(1,1)|\psi\rangle$ of its corresponding eigenstate $|\psi\rangle$ with $|(1,1)\rangle=(b_{1}^{\dagger})^{2}|0\rangle$. Computed for our Hamiltonian $H$ [Eq. 1] at strong non-Hermitian hopping asymmetry $r=4$, it features 5 bands, with bands 2 and 4 from the hybridization of bulk and topological bands. This follows from the single-boson band structure [78, 38, 79, 80, 81, 82] with two symmetrically gapped bulk bands separated by in-gap topological zero modes. The effect of the $U$ density interaction is to induce high $\psi_{11}$ overlap (red) successively from bands 3 to 5, as $U$ is increased.
Figure 4: Persistence of ultrastrong boundary condensation $\bar{P}_{11}$ from interaction-induced caging, even for initial state (4,4) outside a $x_{0}=3$ cage. (a) Snapshots of the time-smoothed ($\Delta t=2$) two-boson correlation probability $\bar{P}_{11}$ (Eq. 6) for (a1) non-interacting and (a2) interacting cases. The bosons still gradually penetrate the L-shaped cage due to the NHSE towards (1,1), and remains trapped in it after that ($t=3,8$). (b1) Eventually, $\bar{P}_{11}(t)$ is ultra-enhanced only for the non-Hermitian interacting case (red), with (b2) negligible enhancement of boundary localization by $U$ in the single-boson case, similar to (b1,b2) of Fig. 3. (c) Time-averaged two-boson probability $\bar{P}_{\text{cage}}$ within the $x_{0}\times x_{0}$ cage as a function of hopping asymmetry $r$. For initial states $(2,2)$ and $(4,4)$ with coincident bosons, the $U=1$ cases exhibit much larger $\bar{P}_{\text{cage}}$ due to interaction-induced caging. But for $(2,4)$, non-coincident initial bosons do not interact appreciably and the effect of interactions is negligible.
@@
To illustrate, with the initial state (1,1), suppressed $\bar{P}_{11}$at $U\approx 3$ corresponds to high $\psi_{11}$ overlap with band 4 [circled in black in Figs. 2(a,b)]. This is because band 4 contains one bulk and one topological boson, which is not consistent with maintaining both bosons at $x_{0}=1$. As such, the same interaction strength of $U\approx 3$ leads to enhanced $\bar{P}_{11}$ (dashed black) for the initial state (1,4), where only one boson overlaps with the topological boundary mode. Likewise, $U\approx 1$ or $U\approx 5$ favors the $\bar{P}_{11}$ bosonic condensation for the bulk initial state (4,4), as circled in light blue, since that is when the $\psi_{11}$ overlap with the bulk bands (3 and 5) is strongest.
@@
2-boson interactions as caging mechanisms.— Interestingly, our density interaction $U$ can also act as a non-local cage when it is situated away from the boundary site $x=1$. In Fig. 3, we demonstrate how this interaction largely confines the evolved state to the left of $x_{0}$ i.e. $x\leq x_{0}$, if the initial state has both bosons in the same region. And, more elaborately, for an initial state with both bosons lying to the right of $x_{0}$, Fig. 4 shows how the interaction will first act as a barrier to the evolved state, but then eventually still trap the particle flux that leaked through it due to NHSE pumping. Overall, this caging mechanism hence further enhances the ultra-strong boson condensation at the boundary. We shall fix the density interaction to be at $x_{0}=3$, and consider initial states $(2,2)$ and $(4,4)$ in Fig. 3 and Fig. 4 respectively.
@@
This caging mechanism is most intuitively represented in the 2D configuration space $(x,x^{\prime})$, where $x$ and $x^{\prime}$ represent the positions of the two bosons [^1] Here $(x,x^{\prime})$ and $(x^{\prime},x)$ refer to the same state due to bosonic statistics [72, 132, 133, 134, 75, 135].. A key observation is that the nonlinear term $\frac{1}{2}Un_{x_{0}}^{2}$ can be interpreted as a non-local L-shaped “cage” [Figs. 3(a) and 4(a)], with potential “walls” [72, 75] of height $\frac{1}{2}U$ across the entire lines $(x,x_{0})$ and $(x_{0},x)$ where only one boson is at $x_{0}$. These walls cross at $(x_{0},x_{0})$, where double bosonic occupancy at $x_{0}$ gives an even higher potential barrier of $2U$.
Fig. 3(a) compares the dynamical evolution of an initial state $(2,2)$ in the (a1) absence and (a2) presence of the density interaction $U$. In Fig. 3(a1) with $U=0$, the smoothed probability cloud $\bar{P}_{xx^{\prime}}(t)$ [Eq. 6] of finding the bosons at $(x,x^{\prime})$ spreads out freely and reflects against the boundaries, even though it is slightly amplified towards the $(1,1)$ corner due to leftwards NHSE pumping. However, in Fig. 3(a2), this spreading is drastically contained within the $3\times 3$ L-shaped cage formed by nonzero $U$ at $x_{0}=3$ [see Supplemental videos]. This caging, which requires both the $U$ interaction and directed amplification towards $(1,1)$, indeed leads to far enhanced boundary boson condensation $\bar{P}_{11}(t)$, as shown in the red curve in Fig. 3(b1).
Interestingly, the caging mechanism protects the $\bar{P}_{11}$ condensation even when the initial state is outside the cage. In Fig. 4(a) with both bosons initially to the right of $x_{0}=3$ at $(4,4)$, their flux $\bar{P}_{xx^{\prime}}(t)$ still gradually enters the $3\times 3$ cage due to the NHSE pumping towards $(1,1)$. Comparing the non-interacting (a1) with the interacting (a2) cases between $t=0$ to $2.2$, the interaction-induced potential walls only slow the diffusion into the cage slightly. However, once the bosons have entered the cage, they are subject to the same trapping mechanism as in Fig. 3, thereby also experiencing eventual ultra-strong condensation at $x=1$ [red in Fig. 4(b1)].
Notably, this caging mechanism is an emergent consequence of few-body density interactions, and cannot be replicated by an effective potential barrier in the single-particle context. Plotted in Figs. 3(b2) and 4(b2) are the evolutions of the densities $\bar{\rho}_{x=1}$ of a single boson in the same 1D SSH chain of Eq. 1, but with an on-site potential $U$ at $x_{0}=3$. Evidently, the $U\neq 0$ cases (orange) do not trap the boson at $x=1$ boundary any more than the $U=0$ cases (blue), with the boundary boson density remaining far lower than the $\bar{P}_{11}(t)$ of the interacting NHSE 2-boson cases [red in Figs. 3(b1) and 4(b1)].
The efficacy of the caging mechanism is further corroborated by plots of $\bar{P}_{\text{cage}}=\sum_{x,x^{\prime}=1}^{x_{0}}\bar{P}_{xx^{\prime}}$, the time-averaged two-boson probability within the cage $x,x^{\prime}\leq x_{0}$. As evident in Figs. 4(c1-c3), $\bar{P}_{\text{cage}}$ increases steadily with the NHSE strength $r$, testimony to the crucial role of non-Hermitian pumping. Comparing Figs. 4(c1) and (c2), we see that the density interaction ($U>0$, orange) significantly enhances $\bar{P}_{\text{cage}}$, particularly when both bosons are already initially inside the cage [Figs.4(c2)]. Saliently, however, the interaction $U$ has a negligible effect if one boson is initially inside the cage and the other outside such that they interact minimally, as in the initial state $(2,4)$ for $x_{0}=3$ [Figs.4(c3)].
Generalizations.— Although we have explicitly considered only two particles, the caging mechanism that leads to ultra-strong boundary localization holds for rather general density interactions with arbitrary numbers of bosons. With $N$ bosons, each Fock state is indexed by a lattice point $\vec{x}=(x_{1},...,x_{N})^{T}$ living in the configuration space $\mathbb{R}^{N}$ [84, 85, 86, 87], and the density takes distinct integer values in the various hyperplanes, depending on the bosonic occupancy. Given a generic interaction that is a multinomial in the densities $n_{1},n_{2},...$ at various sites, a monomial $Un_{x_{1}}n_{x_{2}}...n_{x_{M}}$ gives different energy offset in the corresponding hyperplanes, such as to form the barriers of a “hyper-cage”. For instance, with 4 bosons, an interaction $Un_{x_{0}}n_{x^{\prime}_{0}}$ gives an energy $U$ in the planes where two bosons at fixed at $x_{0}$ and $x^{\prime}_{0}$, $2U$ along lines where three bosons are fixed at either $x_{0}$ or $x^{\prime}_{0}$, etc. Due to the non-locality of the NHSE [22, 23, 14, 24, 25], caging can occur as long as translation invariance is broken in the configuration space, even if the interaction-induced barriers do not completely enclose any region.
@@
Discussion.— We have revealed how the interplay between the NHSE, topology and boson-boson interactions can unexpectedly lead to ultra-strong bosonic condensation. This particle condensation strongly exceeds that of NHSE-pumped free bosons, being also facilitated by the interaction-induced hybridization of topological and bulk states, as well as an emergent non-local caging mechanism. Our findings generalize to higher numbers of bosons as well as generic density interactions and topologies, bringing forth a new approach to trapping and controlling bosons. Experimentally, demonstrations of non-Hermitian systems have been maturely built upon mechanical arrays [88, 89], photonic [90, 91, 62, 92, 29], electrical circuit [24, 55, 93, 81] and quantum circuit platforms [7]. Various platforms can potentially implement the setup for locally interacting bosons to incorporate the $n_{x}^{2}$ term, including optical lattices [94, 95, 96, 97, 98, 99], photonic systems [100, 101, 102, 103, 104, 105, 106, 107, 73], circuit QED [108, 109, 110, 111, 112],various quantum simulators [113, 114, 115, 71, 116, 117, 97, 118, 119, 120, 121, 122, 123, 124, 109, 125, 126, 127, 128, 129, 130, 131] as well as photonic resonator arrays, as further elaborated in [78].
@@
Supplemental Material for “Non-Hermitian ultra-strong bosonic condensation through interaction-induced caging”